OpenStudy (anonymous):

if sin (theta) = -3/4 and cos (theta) >0, find the value of cos (theta)?

OpenStudy (anonymous):

draw a triangle 

OpenStudy (anonymous):

there is a triangle with the opposite over the hypotenuse is \(\frac{3}{4}\) all you need is the third side which you find via pythagoras

OpenStudy (anonymous):

it is \[x=\sqrt{4^2-9^2}=\sqrt{16-9}=\sqrt7\]

OpenStudy (anonymous):



OpenStudy (anonymous):

now you can take any trig ratio you like this makes \[\cos(x)=\frac{\sqrt7}{4}\]